OpenStudy (anonymous):

Evaluate and write the answer in scientific notation: (4x10^5) / (2.5 x 10^8)

OpenStudy (anonymous):

10^5 can be cancelled with 10^8 to yield 10^3 in the denominator. You get 4/(2.5*10^3) = 4/(5000/2) = 8/5000 = 1/625 = ?

OpenStudy (anonymous):

1.6 x 10^ 5/8?